# Toward Universal Decoding of Binary Linear Block Codes via Enhanced Polar Transformations

### v1

@@ V-A The Challenging Case
As previously mentioned, in [14], we exhausted all $40320$ permutation matrix for ${\mathsf{PD}}$ and found that the best sum Bhattacharyya parameter is $0.98291$, which is larger than $0.596991$ obtained by viewing them as independent channels.
To demonstrate the effectiveness of the proposed ${\mathsf{PD}^{+}}$, we first exhaustively explore all $2^{12}\times 40320\approx 1.6\times 10^{8}$ candidate transformations and we identify one that achieves a sum Bhattacharyya parameter of $0.0556$, which is only $6\%$ of the value obtained when treating the channels as independent. The simulation results, presented in Fig. 6, demonstrate that the newly developed ${\mathsf{PD}^{+}}$ achieves MLD performance with only $L=1$ (i.e., SC decoding), whereas ${\mathsf{PD}}$ requires $L=4$ to achieve the same performance, which is only marginally less than the effort required to check all 8 codewords, as in MLD.
@@ V-B Extended BCH Code
We present the results of our proposed ${\mathsf{PD}^{+}}$ for decoding two extended BCH codes: the $(128,54)$ and $(128,106)$ eBCH codes. Due to the prohibitively high complexity of MLD, its simulation is omitted for these codes. In [14], we demonstrated the effectiveness of ${\mathsf{PD}}$ in decoding eBCH codes with $n=64$. However, for $n=128$, due to its immense computational complexity required by an exhaustive search adopted by ${\mathsf{PD}}$, we were unable to identify good instances in [14]. With the introduction of an AI-inspired search algorithm, ${\mathsf{PD}^{+}}$ now enables decoding for codes with $n=128$. The simulation results for the $(128,57)$ eBCH code are presented in Fig. 8, with the complexity of each decoding algorithm detailed in Table I. As shown in the figure and table, the proposed ${\mathsf{PD}^{+}}$ with $L=32$ outperforms OSD with order 2 while visiting significantly fewer candidate sequences. In contrast, GRAND struggles to achieve comparable performance even with 100K visited candidates, primarily due to the drastically increased search space associated with $n-k$.
@@
TABLE II: Frame error rate vs $E_{b}/N_{0}$ for the $(128,106)$ eBCH code.
@@ V-C Extended Golay Code Revisited
We now revisit the $(24,12)$ eGolay code. In [14], we proposed using ${\mathsf{PD}}$ to transform this code into a multi-kernel polar code with dynamic frozen bits, which is then decoded as a polar code. For the newly proposed ${\mathsf{PD}^{+}}$, we begin with a polar code of size $N=32$ using Arikan’s kernel as defined in (4). The kernel is subsequently pruned, and the code is shortened to $n=24$. Simulated annealing is employed to identify good pruning and shortening patterns. Our results, presented in Fig. 10 and Table III, demonstrate that OSD with order 1, ${\mathsf{PD}}$ with a list size of 64, GRAND with 10K candidates, and the proposed ${\mathsf{PD}^{+}}$ with a list size of 8 all achieve near-ML performance. Notably, the proposed ${\mathsf{PD}^{+}}$ achieves this performance with the lowest complexity among the compared methods.
@@ VI Concluding Remark
In this paper, building on the idea of ${\mathsf{PD}}$ presented in [14], we propose a novel universal decoding scheme, ${\mathsf{PD}^{+}}$, for BLBCs. The core concept is to transform the BLBC being decoded into a polar-like code that can be efficiently decoded using existing polar code decoding algorithms. This target polar-like code is constructed by pruning the kernel of Arikan’s polar code and applying code shortening.
@@
Future work includes the following directions: 1) Extending ${\mathsf{PD}^{+}}$ to longer blocklengths: This presents a significant challenge due to the exponential growth of the search space with blocklength; 2) Generalizing ${\mathsf{PD}^{+}}$ to non-binary codes, such as Reed-Solomon codes; 3) Applying more powerful tools from AI and machine learning to solve the local search problem at hand.

### v2

@@ V-A The Challenging Case
As previously mentioned, in [14], we exhausted all $40320$ permutation matrices for ${\mathsf{PD}}$ and found that the best sum Bhattacharyya parameter is $0.979778$, which is larger than $0.596991$ obtained by viewing them as independent channels.
To demonstrate the effectiveness of the proposed ${\mathsf{PD}^{+}}$, we first exhaustively explore all $2^{12}\times 40320\approx 1.6\times 10^{8}$ candidate transformations and we identify one that achieves a sum Bhattacharyya parameter of $0.05536$, which is only $6\%$ of the value obtained when treating the channels as independent. The simulation results, presented in Fig. 6, demonstrate that the newly developed ${\mathsf{PD}^{+}}$ achieves MLD performance with only $L=1$ (i.e., SC decoding), whereas ${\mathsf{PD}}$ requires $L=4$ to achieve the same performance, which is only marginally less than the effort required to check all 8 codewords, as in MLD.
We also provide the simulation result for the proposed ${\mathsf{PD}^{+}}$ using one of the mediocre transformations, which yields a sum Bhattacharyya parameter of 0.08708. A comparison between the two ${\mathsf{PD}^{+}}$ schemes with $L=1$ supports our choice of using the sum of the Bhattacharyya parameters as the optimization criterion. Moreover, we emphasize that for this challenging case, since all 40320 possible permutations for ${\mathsf{PD}}$ were exhausted, the results in Fig. 6 also serve as evidence that incorporating pruning into ${\mathsf{PD}^{+}}$ is crucial for finding improved transformations, as discussed in Remark 5.
@@ V-B Extended BCH Code
We present the results of our proposed ${\mathsf{PD}^{+}}$ for decoding two eBCH codes: the $(128,54)$ and $(128,106)$ eBCH codes. Due to the prohibitively high complexity of MLD, its simulation is omitted for these codes. In [14], we demonstrated the effectiveness of ${\mathsf{PD}}$ in decoding eBCH codes with $n=64$. However, for $n=128$, due to its immense computational complexity required by an exhaustive search adopted by ${\mathsf{PD}}$, we were unable to identify good instances in [14]. With the introduction of an AI-inspired search algorithm, ${\mathsf{PD}^{+}}$ now enables decoding for codes with $n=128$. The simulation results for the $(128,57)$ eBCH code are presented in Fig. 8, with the complexity of each decoding algorithm detailed in Table I. As shown in the figure and table, the proposed ${\mathsf{PD}^{+}}$ with $L=32$ outperforms OSD with order 2 while visiting significantly fewer candidate sequences. In contrast, GRAND struggles to achieve comparable performance even with 100K visited candidates, primarily due to the drastically increased search space associated with $n-k$.
@@
TABLE II: Frame error rate vs $E_{b}/N_{0}$ for the $(128,106)$ eBCH code
@@ V-C Extended Golay Code Revisited
We now revisit the $(24,12)$ eGolay code. In [14], we proposed using ${\mathsf{PD}}$ to transform this code into a multi-kernel polar code with dynamic frozen bits, which is then decoded as a polar code. For the newly proposed ${\mathsf{PD}^{+}}$, we begin with a polar code of size $N=32$ using Arıkan’s kernel as defined in (4). The kernel is subsequently pruned, and the code is shortened to $n=24$. Simulated annealing is employed to identify good pruning and shortening patterns. Our results, presented in Fig. 10 and Table III, demonstrate that OSD with order 1, ${\mathsf{PD}}$ with a list size of 64, GRAND with 10K candidates, and the proposed ${\mathsf{PD}^{+}}$ with a list size of 8 all achieve near-ML performance. Notably, the proposed ${\mathsf{PD}^{+}}$ achieves this performance with the lowest complexity among the compared methods.
@@ V-D Binary Quadratic Residue Code
V-E A Randomly Generated BLBC
Here, we generate an $8\times 16$ generator matrix for a systematic encoder, consisting of an $8\times 8$ identity matrix concatenated with an $8\times 8$ random binary matrix. The elements of the random matrix are independently drawn from $\mathsf{Bernoulli}(0.5)$. The generator matrix is given by
${\scriptsize\mathbf{G}=\left[\begin{array}[]{cccccccccccccccc}1&0&0&0&0&0&0&0&1&1&0&0&1&0&1&0\\ 0&1&0&0&0&0&0&0&0&0&1&0&1&1&0&0\\ 0&0&1&0&0&0&0&0&1&0&1&0&0&1&0&0\\ 0&0&0&1&0&0&0&0&0&0&1&0&1&0&0&1\\ 0&0&0&0&1&0&0&0&0&1&1&0&1&0&0&0\\ 0&0&0&0&0&1&0&0&0&0&1&0&0&1&0&0\\ 0&0&0&0&0&0&1&0&0&1&0&0&1&0&1&0\\ 0&0&0&0&0&0&0&1&1&0&1&0&0&1&1&1\end{array}\right].}$ (25)
We input this $\mathbf{G}$ to the proposed ${\mathsf{PD}^{+}}$, and obtain the transformation $(\mathbf{P},\mathbf{R},\mathcal{S})$ with $\mathbf{P}$ given in (26), $\mathbf{R}$ as illustrated in Fig. 12, and $\mathcal{S}=\emptyset$ (i.e., $\mathbf{S}=\mathbf{I}$).
${\scriptsize\mathbf{P}=\left[\begin{array}[]{cccccccccccccccc}0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&1\\ 0&0&0&1&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&0&0&1&0&0&0&0&0&0\\ 0&0&0&0&0&0&1&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&1&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&1&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&0&1&0&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&1&0&0&0&0&0&0&0\\ 1&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&0&0&0&1&0&0&0&0\end{array}\right].}$ (26)
Figure 12: The pruned kernel and the polarization result. The variables highlighted in red make up the information set.
Simulation results for this randomly generated code, decoded using ${\mathsf{PD}}$ with $\mathbf{P}=\mathbf{I}$ and the proposed ${\mathsf{PD}^{+}}$, are shown in Fig. 13. As illustrated in the figure, applying ${\mathsf{PD}}$ with $\mathbf{P}=\mathbf{I}$ and $L=1$ results in a performance significantly inferior to that of MLD, indicating that this randomly generated BLBC lacks a polar-like structure. In contrast, the performance greatly improves when the proposed ${\mathsf{PD}^{+}}$ is employed, and they closely approach the MLD performance at $L=8$. These results further demonstrate the effectiveness of our proposed transformation in converting a BLBC–even one without inherent polar-like structure–into a form amenable to efficient and near-optimal decoding.
Figure 13: Frame error rate vs $E_{b}/N_{0}$ for the $(16,8)$ code generated by (26).
@@ VI Concluding Remark
In this paper, building on the idea of ${\mathsf{PD}}$ presented in [14], we propose a novel universal decoding scheme, ${\mathsf{PD}^{+}}$, for BLBCs. The core concept is to transform the BLBC being decoded into a polar-like code that can be efficiently decoded using existing polar code decoding algorithms. This target polar-like code is constructed by pruning the kernel of Arıkan’s polar code and applying code shortening.
@@
Future work includes the following directions: 1) Extending ${\mathsf{PD}^{+}}$ to longer blocklengths: This presents a significant challenge due to the exponential growth of the search space with blocklength; 2) Generalizing ${\mathsf{PD}^{+}}$ to non-binary codes, such as Reed–Solomon codes; 3) Applying more powerful tools from AI and machine learning to solve the local search problem at hand.